Sally invested some money at 12% interest. Sally also invested $212
more than 2 times that amount at14%. How much is invested at each rate if Sally receives $945.68 in interest after one year? (Round to two decimal places if necessary.)

Answer by mananth(16949)

Sally invested some money at 12% interest. Assume it is x
Sally also invested $212more than 2 times that amount at14%.
that implies that she invested 2x +212 at 14%
Interest received = $945.68
12%x + 14%(2x+212)= 945.68
multiply the equation by 100
12x+14(2x+212) = 94568
12x+28x+2968 = 94568
40x = 94568-2968
40x=91600
x= 2290
you can continue